# Motion-Zero: Zero-Shot Moving Object Control Framework for Diffusion-Based Video Generation

### v3

@@ 5.1 Implementation Details
Our algorithm is fully implemented during the inference stage, thus it does not require any training. The hyper-parameter $\lambda_{i},\lambda_{o}$ are set to 1, $\lambda_{c}$ is set to 0.05, $\lambda_{s}$ is set to 0.5, $\lambda_{p}$ is set to 0.8. To balance the trade-off between the size of GPU memory consumption and the semantic information retained in the attention map $\mathbf{A}$, we chose a $48\times 48$ size for $\mathbf{A}$ when the output resolution is $384\times 384$. We use DDIM Song et al. (2020) as our sampling method. In the experiment, $T_{1}$ in Fig.2 is set as 10, which means the SC and the STAM are employed during the first 10 timesteps; $T_{2}$ is set to 20, thus the total denoising timestep $T$ is 30. All of our experiments are conducted on a single A100 GPU.
5.2 Comparisons with Others
We evaluate the effectiveness of our method from both qualitative and quantitative perspectives. Two video diffusion baseline methods Sterling (2023); Wang et al. (2023a) are involved to confirm that our approach maintains generative performance. In addition, the SOTA pre-trained model MotionCtrl Wang et al. (2023c) is also involved in experiments to demonstrate the control ability of our proposed method.
Qualitative Results. The results of the qualitative experiments are shown in Fig.3. The base model depicted in the first two rows of the figure is ModelScope. The first row incorporates additional input boxes and utilizes the Motion-Zero framework, while in the second row, an extra prompt is added: moving from left to right. As shown in the first row left, the fish follows a trajectory that strictly adheres to the positions of the boxes. Conversely, in the second row left, the direction of the fish’s movement is not fixed, and the locations where it appears are generated randomly. In the second row right, the bear gradually rotates clockwise and fails to follow the guidance provided by the additional prompts. We can see that the model controlled by the Motion-Zero framework is capable of more effectively governing the trajectory and direction of an object’s movement which cannot be achieved by prompts alone. The base model for the third and fourth rows is ZeroScope. In the third row, both the fish and the bear move according to the conditions set by our boxes, whereas in the fourth row, they remain mostly stationary. This shows that our Motion-Zero model effectively facilitates trajectory control across various baseline models. The fifth row in the figure demonstrates the effects of the pre-trained model MotionCtrl. It can be seen that our results are comparable to those of the pre-trained model in terms of controlling the direction of the movements of the fish. In the case of generating a bear, MotionCtrl only manages to rotate the bear’s body, whereas Motion-Zero can effectively generate the bear’s movement. Note that our method additionally allows for the control of their size. Our advantage also lies in the ability to plug and play with any base model without the need for any training.
Quantitative Results with Automatic Metrics. Followed by LOVEU-TGVE competition Wu et al. (2023b), the CLIP score Hessel et al. (2021) is used to verify text-video consistency (Text Align) and inter-frame consistency (Consistency). The PickScore Kirstain et al. (2023) is used to predict user preferences of our model. We utilized five prompts, with each prompt generating ten videos to serve as test data. The results are shown in Tab.1. From the first four rows of the table, we can discern that employing our Motion-Zero framework to direct the motion trajectory of objects does not compromise the base model’s performance. Instead, the motion subject being constrained by the box results in improved semantic accuracy and temporal consistency. Upon examining the fifth row of the table, it is evident that Motion-Zero achieves an enhancement in performance over pre-trained methodology.
User Study. We recruited 5 voters from social media to evaluate 10 generated videos from the aspect of appearance, temporal consistency, and control capacity, which evaluate whether the position of the moving object is similar to the references. Voters are presented with pairs of videos in a randomized sequence without revealing the creation methods. They are then tasked with assessing each pair and indicating which video they find more favorable. The results are shown in Tab.2. The number in the table denotes the number of Motion-Zero enhanced videos preferred by the voters. v.s. ZeroScope and v.s. ModelScope are the two baseline methods supplemented with guiding prompts, for instance, moving from left to right. v.s. MotionCtrl is a comparison between ZeroScope augmented with our Motion-Zero and the standalone MotionCtrl approach with the same control trajectory. We can see that the voters prefer our methods from all aspects compared with baselines. Moreover, when compared to MotionCtrl, users demonstrated a clear preference for the control capabilities of Motion-Zero. Note that our proposed approach does not need any training. Thus, the baseline model’s inherent limitations resulted in our approach being slightly inferior to MotionCtrl in aspects of consistency.
Text Align Consistency PickScore
ZeroScope 20.05 0.89 18.15
 +Ours 22.04 0.93 20.00
ModelScope 20.30 0.91 18.56
 +Ours 24.74 0.91 19.82
MotionCtrl 20.94 0.92 19.62
Table 1: Automatic metric on baseline methods and SOTA methods. All metrics are such that higher values indicate better performance.
@@ 5.2 Comparisons with Others
v.s. ZeroScope 41/50 20/50 46/50
v.s. ModelScope 38/50 34/50 46/50
v.s. MotionCtrl 25/50 23/50 36/50
Table 2: User study on different motion control methods.
Figure 4: Ablation studies on different components. We use ZeroScope as our baseline model. To demonstrate the coherence of the generated object’s motion, we captured every other frame, resulting in a total of 6 frames. The prompt is A lion is walking on the field. Zoom in for the best view.
5.3 Ablation Study
We conduct a series of ablation experiments to verify the effectiveness of our modules individually.
Impact of INPM. INPM is designed to provide a prior for the initial position of moving objects, facilitating the model to more stably control the position of the object’s movement. As shown in Fig.4, the second row demonstrates the effects of removing this module. We can see that the lion is not entirely within the frame, and due to the forceful control by SC, the lion’s body undergoes significant deformation, while the lion’s head shows no movement at all. For Motion-Zero, the lion’s body is clearly generated, and the lion’s head also moves with the changes in the frame, resulting in the high overall stability of the lion. This shows that the INPM module plays a significant role in stably controlling the position of objects and generating stable videos.
Impact of SC. In this section, we validate the importance of SC. SC primarily provides auxiliary support for control over spatial positioning. As shown in the third line in Fig.4, without spatial constraints, the generated motion of the lion appears mostly outside the box. The SC is the most important module for our Motion-Zero controlling method. The reason is that SC is the only gradient-based modification to the baseline compared with other modules.
Impact of STAM. The role of the STAM is to ensure the temporal consistency of the moving object, which is to maintain the continuity of its motion as it moves. As illustrated in Fig.4, we present the effects of omitting the STAM module in the last row. It can be noted that during the movement, the lion’s legs undergo deformation and become blurred. In contrast, the first row demonstrates that the lion’s leg movements are well-generated with strong continuity, highlighting the significance of our STAM. More details are provided in the supplementary materials.
6 Discussion, Limitation and Future Works
Due to the absence of the need for additional training, our method can be applied to all pre-trained video diffusion models. However, on the other hand, its generative performance is entirely dependent on the base model. Existing video diffusion models sometimes generate relatively unstable objects, such as deformation in moving objects, holes in the background, and various issues like the inability to precisely control the attributes of generated objects with text.
Our method can currently effectively control the movement of objects in videos, but the trajectory of the movement is entirely controlled by the user, lacking semantic interaction with the video. For example, in a forest scene, if a user provides the first frame as a reference, they can use a prompt like a little rabbit goes to the stream to drink water to control a rabbit to navigate and avoid obstacles in the forest automatically. This kind of interaction between object movement and the background, as well as narrative-like semantic control, will be the direction of our future work.
7 Conclusion
@@
In this paper, we proposed a framework Motion-Zero for controlling object motion that can be applied to any video diffusion model. Unlike previous methods that require extensive training, our method enables arbitrary motion control in a zero-shot setting. We find that the initial noise plays a significant role in object motion, leading us to introduce an initial noise prior module to achieve stable control effects. We discovered that the values in the cross-attention map significantly influence the position and appearance of the object corresponding to the prompt. Therefore, spatial constraints are designed to optimize the position of moving objects and ensure spatial consistency. Furthermore, to ensure the consistency of object motion, a shift temporal attention mechanism is developed to maintain the temporal coherence of moving objects. Extensive experiments substantiate the efficacy and generalization of our approach.

### v4

@@ Introduction
Our algorithm is fully implemented during the inference stage, thus it does not require any training. The hyper-parameter $\lambda_{i},\lambda_{o}$ are set to 1, $\lambda_{c}$ is set to 0.05, $\lambda_{s}$ is set to 0.5, $\lambda_{p}$ is set to 0.8. To balance the trade-off between the size of GPU memory consumption and the semantic information retained in the attention map $\mathbf{A}$, we choose a $48\times 48$ size for $\mathbf{A}$ when the output resolution is $384\times 384$. We use DDIM (Song, Meng, and Ermon 2020) as our sampling method. In the experiment, $T_{1}$ in Fig.2 is set as 10, which means the SC and the STAM are employed during the first 10 timesteps; $T_{2}$ is set to 20, thus the total denoising timestep $T$ is 30. All of our experiments are conducted on a single NVIDIA A100 GPU.
Comparisons with SOTA Methods
Qualitative Analysis.
The results of the qualitative experiments with simple and complex trajectories are shown in Fig.3 and Fig.4, respectively. Our proposed Motion-Zero (+Ours) applied on other baselines (ModelScope and ZeroScope) can greatly increase the controllability of the objects’ motion trajectories. In Fig.3, baselines (+Ours) refers to applying MotionZero to the baseline methods, while original baselines (+Prompt) have an extra prompt added to indicate the motion of objects: moving from left to right. As shown in Fig. 3, Motion-Zero correctly guides the motion of fishes by following the specific trajectories indicated by the red boxes. In contrast, both original ModelScope (+Prompt) and ZeroScope (+Prompt) fail to control the objects following the expected trajectories. Fig.3 (c) demonstrates the generated results of TrailBlazer. It shows that the fish generated by TrailBlazer does not strictly follow the red bounding boxes. Fig.3 (d) demonstrates the results of Peekaboo. It is observed that the object detaches from the box in the last few frames, and the movement direction of the fish on the left does not align with the direction of the box.
Fig.4 indicates the comparison between ours, TrailBlazer, and Peekaboo in complex trajectories setting. Our method effectively controls the motion of objects for any trajectory, e.g. a flying back rocket. On the contrary, TrailBlazer exhibits cases where moving objects are coupled with the background motion, e.g. the penguin is relatively stationary. Furthermore, the rocket and the rabbit are losing control and the motion becomes irrelevant with the bounding boxes. The penguins generated by Peekaboo lack dynamism, maintaining a single pose while only the scene moves. It fails to properly generate rockets, which might be due to the unusual nature of rotating and moving rockets. Regarding the rabbit scene, the content generated by Peekaboo exhibited frame skipping.
Figure 3: Quality comparison results on different methods. We take one frame from every three frames. The input prompt: A fish is swimming in the sea. We employed ModelScope (a) and ZeroScope (b) as our baseline models and compared the effect of incorporating additional prompts with the integration of our Motion-Zero. In addition, we conducted a comparative analysis with TrailBlazer and Peekaboo.
Figure 4: Quality comparison results with complex trajectories. We take one frame from every three frames. The input prompt of the first row: A penguin standing on an iceberg. The second row: A rocket launching into space from a launchpad. The third row: A rabbit burrowing downwards into its warren. Zoom in for the best view.
Quantitative Analysis.
Following LOVEU-TGVE competition (Wu et al. 2023b), we use the CLIP score (Hessel et al. 2021) to verify text-video consistency (Text Align) and inter-frame consistency (Consistency). The PickScore (Kirstain et al. 2023) is employed to predict user preferences of our model. To further evaluate the control capability of our model, we employ metrics including mIoU, AP50, Cov., and CD. mIoU stands for mean Intersection-over-Union of the detected bounding boxes and the input box on the generated video. This metric is primarily used to assess whether the model can effectively control the position and size of moving objects. We use the OWLViT-large detector (Minderer et al. 2022) to detect object boxes in the generated video. AP50 refers to average precision@50%, which is used to determine if the overlap between the detected boxes and the user-provided boxes is greater than 50%. Cov. is used to assess the probability of the detector detecting moving objects. CD represents the distance between the center of the moving object in the generated video and the center of the user-provided box, reflecting whether the generated center of the moving object is consistent with the user input.
Align Cons. Pick. mIoU AP50 Cov. CD($\downarrow$)
ZeroScope 20.31 0.88 18.98 - - - -
 +Ours 21.96 0.94 19.89 0.54 0.64 0.96 0.07
ModelScope 20.55 0.91 18.70 - - - -
 +Ours 23.54 0.90 19.56 0.54 0.64 0.88 0.08
TrailBlazer 21.53 0.92 19.64 0.52 0.64 0.91 0.14
Peekaboo 20.04 0.83 18.70 0.43 0.45 0.89 0.11
TrailBlazer (c.) 23.34 0.92 18.81 0.50 0.61 0.91 0.13
Peekaboo (c.) 20.89 0.84 18.53 0.41 0.42 0.88 0.11
Ours (c.) 22.40 0.95 19.19 0.55 0.67 0.97 0.07
Table 1: Automatic metric on baseline methods and SOTA methods. The left half indicates the quality of the generation, while the right half demonstrates the control capability of the model. All metrics expect CD to be such that higher values($\uparrow$) indicate better performance. (c.) means the method is tested in complex trajectories. Align means Text Align, Cons. means Consistency and Pick. means PickScore.
As shown in Tab.1, the left part indicates the generation quality, and it shows that our proposed Motion-Zero framework does not compromise the baseline model’s performance but increases video generation quality instead. This demonstrates that our method exploits and preserves the model’s generation capability even with the zero-shot setting. In addition, the motion subject constrained by the box results in improved semantic accuracy and temporal consistency. This indicates that our proposed Motion-Zero can get a better generation quality score due to the effectiveness of consistency between the prompt and motion.
For the left in Tab.1, our method outperforms Trailblazer in terms of control performance for both simple and complex trajectories and surpasses Peekaboo in every aspect. Trailblazer performs better on the Align metric because it manually modifies cross-attention rather than deriving it through loss. However, this diminishes its performance on other metrics. Through mIoU and AP50 metrics, we can see that the size and position of the moving objects generated by our method have a closer match for the boxes provided by the users. This confirms that $L_{o}$ and $L_{i}$ indeed ensure that the generated objects remain within the boxes given by the users. Higher Cov. score demonstrates that the objects generated by our method are clearer. This proves that our INPM and STAM modules enable smoother motion of objects, while $L_{s}$ maintains the consistency of the objects. At the same time, CD indicates that the centers of the objects generated by our method are closer to the centers of the boxes provided. This reflects that $L_{c}$ can position the generated objects at the center of the box.
Align Cons. Pick. mIoU AP50 Cov. CD($\downarrow$)
w/o INPM 20.30 0.93 19.29 0.40 0.41 0.86 0.09
w/o SC 21.27 0.92 19.10 0.18 0.17 0.71 0.22
w/o STAM 20.64 0.93 19.27 0.43 0.47 0.83 0.13
Ours 21.96 0.94 19.89 0.54 0.64 0.96 0.07
Table 2: The table of ablation study on different modules.
Ablation Study
Quantitative Analysis. We conduct ablation experiments on different components of Motion-Zero, as shown in Tab.2. It indicates that removing individual modules leads to a significant decrease in the model’s performance. We observe that when SC is removed, the scores of Align do not decrease significantly. However, when STAM or INPM is removed, the scores of Align decrease noticeably, demonstrating the significant role of the STAM and INPM modules in improving the consistency between text and generated content. The SC has the greatest impact on location control as the scores drop from 0.54 to 0.18 in mIoU and from 0.64 to 0.17 in AP50. Without the SC module, there is no guarantee for objects’ position and spatial consistency between frames. Additionally, the absence of STAM and INPM also results in inferior performance in temporal consistency for moving objects as the Cov. drops and it becomes difficult for the model to generate expected positioned objects as mIoU and AP50 decrease. Overall, the three modules are all important for video generation and the STAM and INPM have more impact on generation quality.
Visualization of Attention Maps.
Figure 5: Attention maps with different components. Prompt: A seal walking on the ice.
We visualize the attention maps in the cross-attention block, as shown in Fig.5. The second row displays the cross-attention map at step 0 under random initial noise conditions. It is observed that the attention values are random. As shown in the third row, when we introduce INPM, we can see that the areas with high attention response at step 0 are already approximately located at the positions of the user’s box. This reflects that INPM can provide strong prior information for the position of moving objects. After optimization by SC, in the fourth and fifth rows, we can see that the areas with high response in the attention map are accurately located at the box position and are close to the position of the generated moving object. This demonstrates that SC can enforce high attention values inside the box while keeping the low response values outside the box.
User Study
@@
ZeroScope 2.97 2.80 2.27
ModelScope 2.57 2.55 2.23
Trailblazer 3.32 3.32 3.40
Peekaboo 3.96 3.84 3.55
Ours 4.75 4.57 4.67
Table 3: The table of user study.
We randomly pick 3 videos from 10 videos generated by each model, 15 in total to be used in our user study. Every video has 3 dimensions of evaluation rating: Appearance, Consistency, and Control, and the scores 1 to 5 indicate appearance, consistency, and control capabilities from low to high. We randomly choose 60 users from our university through an incentivized questionnaire to evaluate each dimension of the 15 videos. At the beginning of our questionnaire, we provide users with a detailed explanation of the task and the scoring guidelines. Only after confirming that the user fully understands these instructions do they proceed with the questionnaire. The results are shown in Tab.3. For baseline models, additional prompts are used to control the movements. The voters prefer our methods from all aspects compared with baselines, Trailblazer, and Peekaboo. To further validate the effectiveness of our user study, we use the Cronbach’s $\alpha$ coefficient to assess the internal consistency (reliability) of the questionnaire. The Cronbach’s $\alpha$ coefficient for our questionnaire is 0.901, indicating a high level of reliability. We use the Friedman Test to confirm that our results are statistically significant. The results indicate significant differences among the five methods (Friedman Test, $X^{2}$ = 1534.29, $p$ < 0.001).
Conclusion
@@
In this paper, we proposed a novel zero-shot framework, Motion-Zero, for arbitrary object motion trajectory control that can be applied to various video diffusion models. Unlike previous methods require extensive training, our method enabled motion-control video generation without any fine-tuning of the baseline model. Our Initial Noise Prior Module was proposed to acquire prior initial noise for a high-quality generation. Moreover, Spatial Constraints and Shift Temporal Attention Mechanism respectively exploited the cross-attention dimension and temporal-attention dimension to obtain spatial control and temporal consistency. Extensive experiments demonstrated the efficacy and generalization of our proposed approach.
Acknowledgments
This work is supported by the National Natural Science Foundation of China (62472178, 62376244), Fundamental Research Funds for the Central Universities, and Shanghai Urban Digital Transformation Special Fund Project (202301027). This work is also sponsored by Natural Science Foundation of Chongqing, China (CSTB2022NSCQ-MSX0552) and the Open Projects Program of State Key Laboratory of Multimodal Artificial Intelligence Systems (No.MAIS2024111).